# Restructuring Tractable Probabilistic Circuits

### v1

@@ 5 PC DEPTH REDUCTION
1:procedure BalancedVtree($V,\mathbf{S}_{l}=\emptyset,\mathbf{S}_{r}=\emptyset$)
@@
3:         return leaf($V$; $\mathbf{S}_{l}\cup\mathbf{S}_{r})$
@@
11:     $V_{l}^{\prime}\leftarrow\textsc{BalancedVtree}(V_{[v\mapsto l]},\mathbf{S}_{l},\{Z_{v}\})$
12:     $V_{r}^{\prime}\leftarrow\textsc{BalancedVtree}(V_{r},\{Z_{v}\},\mathbf{S}_{r})$
13:     return $\textsc{join}(V_{l}^{\prime},V_{r}^{\prime};\mathbf{S}_{l}\cup\mathbf{S}_{r})$
@@
Depth reduction of a probabilistic circuit refers to the construction of an equivalent circuit with reduced depth, e.g. to a depth logarithmic in the number of variables. A depth reduction algorithm for general circuits is known (Valiant et al.,, 1983; Raz and Yehudayoff,, 2008; Yin and Zhao,, 2024) but does not take advantage of structuredness. We show how to reduce a structured-decomposable circuit to an equivalent log-depth circuit by restructuring. The key step is to identify a log-depth vtree such that restructuring to that vtree using Algorithm 1 (and some valid choice of labels) results in at most a polynomial increase in size.
@@
Theorem 5.1 (Depth Reduction Vtree).
Given any vtree $V$, Algorithm 4 returns a vtree $W$ of depth $O(\log|V|)$ with a valid labelling of cardinality $3$.
@@
The depth reduction to $O(\log|V|)$ is achieved as the algorithm increases the depth by one in each recursive call, but reduces the vtree size by a multiplicative factor. The validity condition holds due to the separation into connected components (the labels can also be obtained from Algorithm 2). The value of $M^{\prime}$ follows by noting that $\mathbf{S}_{l}$ and $\mathbf{S}_{r}$ are either empty or singleton sets, and that the algorithm produces $\mathbf{C}_{w}=\mathbf{S}_{l}\cup\mathbf{S}_{r}$, $\mathbf{C}_{l}=\mathbf{S}_{l}\cup\{Z_{v(w)}\}$, and $\mathbf{C}_{r}=\{Z_{v(w)}\}\cup\mathbf{S}_{l}$ where $Z_{v(w)}$ for each inner node $w\in W$. ∎
@@
Firstly, the depth-reduced PC retains structuredness, which is not guaranteed by the existing depth-reduction algorithms. Secondly, exploiting structuredness and tracking the hidden state size enables a more fine-grained analysis of the size of the depth-reduced circuit. Since the size of the original circuit is $O(nh^{2})$, using the known cubic bound on the size of the depth-reduced circuit (Raz and Yehudayoff,, 2008) gives $O(n^{3}h^{6})$. However, by Theorem 5.1, we see that $M^{\prime}\!=\!\max_{w\in W}|\mathbf{C}_{l},\mathbf{C}_{r},\mathbf{C}_{W}|\!\leq\!3$ and so by Theorem 3.7 we immediately obtain a much tighter bound of $O(nh^{3})$ for the size of the resulting circuit.
@@ Acknowledgements
This work was funded in part by the DARPA ANSR program under award FA8750-23-2-0004, the DARPA PTG Program under award HR00112220005, and NSF grant #IIS-1943641. This work was done in part while BW, MA and GVdB were visiting the Simons Institute for the Theory of Computing.
@@ Appendix A ADDITIONAL PROOFS
Suppose that $\mathscr{A}$ is a structured decomposable and smooth PC respecting vtree $V$. Write $\text{prod}(v)$ and $\text{sum}(v)$ for the set of product and sum nodes with scope $\mathbf{X}_{v}$. The augmented PC $\mathscr{A}_{\text{aug}}$ is decomposable as if leaves with scope $\{Z_{v}\}$ were contained in (the subcircuits rooted at) two different children $t_{1},t_{2}$ of a product node, then their parents (nodes in $\text{prod}(v)$) would be contained in $t_{1},t_{2}$, which is a contradiction of decomposability of $\mathscr{A}$. It is also smooth as for any sum node, if one sum node contains some leaf with scope $\{Z_{v}\}$, then it contains some node in $\text{sum}(v)$, hence by smoothness of $\mathscr{A}$ all sum nodes contain some node in $\text{sum}(v)$ and thus some leaf with scope $\{Z_{v}\}$.
Consider the standard marginalization algorithm for PCs (Darwiche,, 2003; Choi et al.,, 2020), where one replaces each leaf whose scope is contained within the variables being marginalized out with the constant $1$. This correctly marginalizes the function represented by the PC if the PC is decomposable and smooth. If we marginalize over all newly introduced latents $\mathbf{Z}$, it is immediate that the resulting PC represents the same function as $\mathscr{A}$. ∎
@@
In Section 3.1 we described a Bayesian network $p_{G}=p^{*}$ with the required graph. It remains to show that this network represents the same distribution as $\mathscr{A}$. We will do this by showing that the Bayesian network has the same distribution as the augmented PC, i.e. $p_{G}(\mathbf{X},\mathbf{Z})=p_{\mathscr{A}_{\text{aug}}}(\mathbf{X},\bm{Z})$.
@@
$p_{\mathscr{A}_{\text{aug}},T}(\mathbf{X},\mathbf{Z}):=\prod_{(t_{i},t_{j})\in\textsc{sumedges}(T)}w_{t_{i},t_{j}}\prod_{t\in\textsc{leaves}_{\mathbf{X}}(T)}f_{t}(X_{\text{sc}(t)})\prod_{t\in\textsc{leaves}_{\mathbf{Z}}(T)}f_{t}(Z_{\text{sc}(t)})$ (1)
where $\textsc{sumedges}(T)$ denotes the set of outgoing edges from sum nodes in $T$, and $\textsc{leaves}_{\mathbf{X}}(T),\textsc{leaves}_{\mathbf{Z}}(T)$ denote the set of leaf nodes in $T$ with scope corresponding to a variable in $\mathbf{X},\mathbf{Z}$ respectively. The distribution of the augmented PC is then in fact given by the sum of these functions over all induced trees:
@@
$p_{\mathscr{A}_{\text{aug}}}(\mathbf{X},\mathbf{Z})=\sum_{T\in\mathcal{T}}p_{\mathscr{A}_{\text{aug}},T}(\bm{X},\bm{Z})$.
Now, let $\text{path}(v,i,j)$ be a predicate indicating whether there is a path between $t_{p,j}$ and $t_{v,i}$ (where we use $p$ to denote the parent vtree node of $v$, and as before e.g. $t_{v,i}$ indicates the product node with scope $\mathbf{X}_{v}$ and corresponding to $Z_{v}=i$). We will consider two cases depending on the value of the latents. Specifically, we will say that an assignment $\bm{z}$ is consistent if $\text{path}(v,z_{v},z_{p})$ holds for all non-root inner nodes in the vtree, and inconsistent otherwise.
If an assignment $\bm{z}$ is inconsistent, then for any assignment $\mathbf{x}$ of the observed variables, we have that $p_{G}(\mathbf{x},\mathbf{z})=0$ by definition of the Bayesian network distribution. Now consider any induced subtree $T\in\mathcal{T}$. Each $T$ must contain one product node for every variable scope. In particular, $T$ must contain some product node $t_{v,i}$ such that $z_{v}\neq i$ (otherwise, since $\mathbf{z}$ is inconsistent, a (connected) tree would be impossible). We then have $p_{\mathscr{A}_{\text{aug}},T}(\bm{x},\bm{z})=0$ for all $\bm{x}$, as Equation 1 then contains a leaf function $f_{t}(z_{v})=\mathds{1}_{z_{v}=i}=0$. Thus $p_{\mathscr{A}_{\text{aug}}}(\mathbf{x},\mathbf{z})=\sum_{T\in\mathcal{T}}p_{\mathscr{A}_{\text{aug}},T}(\bm{x},\bm{z})=0$ for any $\mathbf{x}$.
If an assignment $\mathbf{z}$ is consistent, note that, by our assumption of alternating sums and products, there can be exactly one path from $t_{p,j}$ to $t_{v,j}$, as $t_{p,j}$ has a unique sum node child with scope containing $\mathbf{X}_{v}$, and this sum node must immediately have $t_{v,j}$ as a child. Thus there is exactly one induced tree $T$ containing $t_{v,z_{v}}$ for all (non-root) inner vtree nodes $v$. Further, examining the definition of the Bayesian network distribution $p_{G}(\bm{X},\bm{z})$, this exactly matches the definition of $p_{\mathscr{A}_{\text{aug}},T}(\bm{X},\bm{z})$: each sum node edge weight in the tree corresponds to a sum node edge weight along a path from some $t_{p,z_{p}}$ to $t_{v,z_{v}}$ and thus the CPT of $Z_{v}$ given $Z_{p}$ (the root sum node edge weight corresponds to the CPT for $Z_{\textsc{root}(V)}$), and each leaf node distribution for observed variables corresponds to the CPT for that variable given its parent.
@@
We first show that $\mathbf{C}_{w}$ satisfies the following properties:
@@
    $X_{a:b}=\bigcup_{Z_{i}\in\mathbf{C}_{a:b}}\textsc{Leaves}(Z_{i})$ is a disjoint union.
@@
    If $a\!\leq\!c\!\leq d\!\leq b$, then for $Z\!\in\!\mathbf{C}_{c:d}$, there exists $Z^{\prime}\in\mathbf{C}_{a:b}$ such that $Z^{\prime}$ is an ancestor of $Z$ in $G_{\mathscr{A}}$.
Property 1 follows from the proof of correctness of the segment tree querying algorithm. Property 2 follows from Property 1 together with the key observation that we can compute $\mathbf{C}_{c:d}$ via $\bigcup_{Z_{i}\in\mathbf{C}_{a:b}}\textsc{SegmentCover}(Z_{i},\mathbf{X}_{c:d}\cap\textsc{Leaves}(Z_{i}))$. Let $w=a\!:\!b$ be a node in $W$ with children $l=a\!:\!c$ and $r=c+1\!:\!b$; it follows from Property 1 that $\mathbf{C}_{a:b}$ covers $\mathbf{X}_{a:b}$ and $\mathbf{C}_{a:c}$ blocks all paths from $\mathbf{X}_{a:c}$ to $\mathbf{C}_{c+1:b}$; it follows from Property 2 that $\mathbf{C}_{a:c}$ blocks all paths from $\mathbf{X}_{a:c}$ to $\mathbf{X}_{a:b}$. Hence we conclude that $\mathbf{C}_{w}$ is a valid labelling. A minor catch is that $\mathbf{C}_{w}$ may contain variables in $\mathbf{X}$, but we can replace them by their parent in $G_{\mathscr{A}}$ without affecting the validity of $\mathbf{C}_{w}$. ∎
@@
By Theorem 5.1, given a structured PC $\mathbf{X}$ over $n$ variables with hidden state size $h$ and respecting vtree $V$, we can generate a vtree $W$ of depth $O(\log n)$ and with labelling cardinality $M^{\prime}=3$. Thus, by Theorem 3.7 we can construct a PC representing the same function and respecting vtree $W$ of size $O(nh^{3})$. The depth of the PC is then also $O(\log n)$ as we have assumed alternating sum and product nodes, so the depth of the circuit is at most double that of the vtree. ∎
@@ Appendix B Computing Minimum D-separators
procedure MinimumSeparator($Z$, $\mathbf{A}$, $\mathbf{B}$)
     if $\mathbf{A}\!=\!\varnothing$ and $\mathbf{B}\!=\!\varnothing$ then
@@
           $Z_{i},\mathbf{A}\!\cap\!\textsc{Leaves}(Z_{i}),\mathbf{B}\!\cap\!\textsc{Leaves}(Z_{i}))$
@@
     $\mathbf{C}_{\mathbf{A}}\leftarrow\textsc{Min}({\bigcup}_{i}\mathbf{C}_{i}\cup\{Z\},{\bigcup}_{i}\mathbf{C}_{i,\mathbf{A}})$
     $\mathbf{C}_{\mathbf{B}}\leftarrow\textsc{Min}({\bigcup}_{i}\mathbf{C}_{i}\cup\{Z\},{\bigcup}_{i}\mathbf{C}_{i,\mathbf{B}})$
     $\mathbf{C}\leftarrow\textsc{Min}(\mathbf{C_{A}},\mathbf{C_{B}})$
     return $\mathbf{C_{A}},\mathbf{C_{B}},\mathbf{C}$
@@
Let $G$ be a tree-shaped Bayesian network rooted at $Z$; in particular, assume that the leaves of $G$ $\subseteq\mathbf{X}\!\cup\!\mathbf{Z}$ and the internal nodes of $G$ $\subseteq\mathbf{Z}$ (e.g. Figure 1(b)). Then, we want to prove that Algorithm 5 computes a minimum d-separator $\mathbf{C}\subseteq G$ for $\mathbf{A},\mathbf{B}\subseteq\mathbf{X}$.
As shown in Algorithm 5, given tree-shaped Bayesian network $G$ rooted at $Z$, the procedure MinimumSepartor computes three sets of latent variables $\mathbf{C_{A}}$, $\mathbf{C_{B}}$ and $\mathbf{C}$. Specifically, we shall prove the following properties:
@@
    $\mathbf{\mathbf{C_{A}}}$ is a minimum d-separator between $\mathbf{A}$ and $\mathbf{B}$ that also blocks all paths from $\mathbf{A}$ to $Z$.
@@
    $\mathbf{\mathbf{C_{B}}}$ is a minimum d-separator between $\mathbf{A}$ and $\mathbf{B}$ that also blocks all paths from $\mathbf{B}$ to $Z$.
@@
    Either $\mathbf{C_{A}}$ or $\mathbf{C_{B}}$ is a minimum d-separator between $\mathbf{A}$ and $\mathbf{B}$ in $G$ (rooted at $Z$); hence $\mathbf{C}$ is a minimum d-separator between $\mathbf{A}$ and $\mathbf{B}$ in $G$.
@@
We prove Property 1 (and Property 2) by a bottom-up induction on $G$. First of all it is easy to verify that the three base cases, i.e. $\mathbf{A}=\varnothing$ and $\mathbf{B}=\varnothing$, $\mathbf{A}=\varnothing$ and $\mathbf{B}=\varnothing$, are correct.
We now prove Property 1 (the proof for Property 2 is symmetric) by induction; first of all it is clear that both ${\bigcup}_{i}\mathbf{C}_{i}\cup\{Z\}$ and ${\bigcup}_{i}\mathbf{C}_{i,\mathbf{A}}$ form d-separators between $\mathbf{A}$ and $\mathbf{B}$, and it remains to show that the minimum of these two is a minimum d-separator between $\mathbf{A}$ and $\mathbf{B}$ that blocks all paths from $\mathbf{A}$ to $Z$. Suppose, towards a contradiction, let $\mathbf{C_{A}}^{\prime}$ be such a d-separator of size $<\textsc{Min}(|{\bigcup}_{i}\mathbf{C}_{i}\cup\{Z\}|,|{\bigcup}_{i}\mathbf{C}_{i,\mathbf{A}}|)$. There are two cases:
  • If $\mathbf{C_{A}}^{\prime}$ contains $Z$: let $G_{i}$ be the subtree rooted at $Z_{i}$ and set ${\mathbf{C}_{i}}^{\prime}=\mathbf{C_{A}}^{\prime}\cap G_{i}$. It is immediate that $\mathbf{C_{i}}^{\prime}$ is a d-separator between $\mathbf{A}$ and $\mathbf{B}$ in $G_{i}$ and by assumption, there exists at least one one $i$ such that $|{\mathbf{C}_{i}}^{\prime}|<|\mathbf{C}_{i}|$; contradicting the induction hypothesis.
  • If $\mathbf{C_{A}}^{\prime}$ does not contain $Z$: let $G_{i}$ be the subtree rooted at $Z_{i}$ and set ${\mathbf{C}_{i,\mathbf{A}}}^{\prime}=\mathbf{C_{A}}^{\prime}\cap G_{i}$. Similarly, it is not hard to see that ${\mathbf{C}_{i,\mathbf{A}}}^{\prime}$ is a d-separator between $\mathbf{A}$ and $\mathbf{B}$ in $G_{i}$ that blocks all paths from $\mathbf{A}\cap G_{i}$ to $Z_{i}$. By assumption, there exists at least one $i$ such that $|{\mathbf{C}_{i,\mathbf{A}}}^{\prime}|<|{\mathbf{C}_{i,\mathbf{A}}}|$; contradicting the induction hypothesis.

### v2

@@ 5 PC DEPTH REDUCTION
1:procedure BalancedVtree($V,\bm{S}_{l}=\emptyset,\bm{S}_{r}=\emptyset$)
@@
3:         return leaf($V$; $\bm{S}_{l}\cup\bm{S}_{r})$
@@
11:     $V_{l}^{\prime}\leftarrow\textsc{BalancedVtree}(V_{[v\mapsto l]},\bm{S}_{l},\{Z_{v}\})$
12:     $V_{r}^{\prime}\leftarrow\textsc{BalancedVtree}(V_{r},\{Z_{v}\},\bm{S}_{r})$
13:     return $\textsc{join}(V_{l}^{\prime},V_{r}^{\prime};\bm{S}_{l}\cup\bm{S}_{r})$
@@
Depth reduction of a probabilistic circuit refers to the construction of an equivalent circuit with reduced depth, e.g. to a depth logarithmic in the number of variables. A depth reduction algorithm for general circuits is known (Valiant et al.,, 1983; Raz and Yehudayoff,, 2008; Yin and Zhao,, 2024) but does not take advantage of structuredness. We show how to reduce a structured-decomposable circuit to an equivalent log-depth circuit by restructuring. Unlike in multiplication, we are not concerned with restructuring to a particular vtree, but rather any log-depth vtree. The key step is thus to identify a log-depth vtree such that restructuring to that vtree using Algorithm 1 (and some valid choice of labels) results in at most a polynomial increase in size.
@@
Theorem 5.1 (Depth Reduction).
Given any vtree $V$, Algorithm 4 returns a vtree $W$ of depth $O(\log|V|)$ with a valid labelling of cardinality $M^{\prime}$ at most $3$.
@@
The depth reduction to $O(\log|V|)$ is achieved as the algorithm increases the depth by one in each recursive call, but reduces the vtree size by a multiplicative factor. The validity condition holds due to the separation into connected components (the labels can also be obtained from Algorithm 2). The value of $M^{\prime}$ follows by noting that $\bm{S}_{l}$ and $\bm{S}_{r}$ are either empty or singleton sets, and that the algorithm produces $\bm{C}_{w}\!=\!\bm{S}_{l}\cup\bm{S}_{r}$, $\bm{C}_{l}\!=\!\bm{S}_{l}\cup\{Z_{v(w)}\}$, and $\bm{C}_{r}\!=\!\{Z_{v(w)}\}\cup\bm{S}_{l}$ where $Z_{v(w)}$ for each inner node $w\!\in\!W$. ∎
@@
Firstly, the depth-reduced PC retains structuredness, which is not guaranteed by the existing depth-reduction algorithms. Secondly, exploiting structuredness and tracking the hidden state size enables a more fine-grained analysis of the size of the depth-reduced circuit. Since the size of the original circuit is $O(nh^{2})$, using the known cubic bound on the size of the depth-reduced circuit (Raz and Yehudayoff,, 2008) gives $O(n^{3}h^{6})$. However, by Theorem 5.1, we see that $M^{\prime}\!=\!\max_{w\in W}|\bm{C}_{l},\bm{C}_{r},\bm{C}_{W}|\!\leq\!3$ and so by Theorem 3.8 we immediately obtain a much tighter bound of $O(nh^{3})$ for the size of the resulting circuit.
@@ 7 DISCUSSION
Our work takes a step towards understanding when and how the space of structured circuits – hierarchical tensor factorizations (Loconte et al., 2024a, ) – can be “connected” through tractable transformations. Many interesting theoretical problems remain open: for instance, the optimality of our algorithm, and circuit lower bounds for particular structures. We also envisage that our restructuring algorithm will enable new methodology both within the PC community and broadly in structured representations of high-dimensional tensors and probability distributions.
@@ Acknowledgements
This work was funded in part by the DARPA ANSR program under award FA8750-23-2-0004, the DARPA CODORD program under award HR00112590089, NSF grant #IIS-1943641, and gifts from Adobe Research, Cisco Research, and Amazon. This work was done in part while BW, MA and GVdB were visiting the Simons Institute for the Theory of Computing.
@@ Appendix A ADDITIONAL PROOFS
Suppose that $\mathscr{A}$ is a structured decomposable and smooth PC respecting vtree $V$. Write $\text{prod}(v)$ and $\text{sum}(v)$ for the set of product and sum nodes with scope $\bm{X}_{v}$. The augmented PC $\mathscr{A}_{\text{aug}}$ is decomposable as if leaves with scope $\{Z_{v}\}$ were contained in (the subcircuits rooted at) two different children $t_{1},t_{2}$ of a product node, then their parents (nodes in $\text{prod}(v)$) would be contained in $t_{1},t_{2}$, which is a contradiction of decomposability of $\mathscr{A}$. It is also smooth as for any sum node, if one sum node contains some leaf with scope $\{Z_{v}\}$, then it contains some node in $\text{sum}(v)$, hence by smoothness of $\mathscr{A}$ all sum nodes contain some node in $\text{sum}(v)$ and thus some leaf with scope $\{Z_{v}\}$.
Consider the standard marginalization algorithm for PCs (Darwiche,, 2003; Choi et al.,, 2020), where one replaces each leaf whose scope is contained within the variables being marginalized out with the constant $1$. This correctly marginalizes the function represented by the PC if the PC is decomposable and smooth. If we marginalize over all newly introduced latents $\bm{Z}$, it is immediate that the resulting PC represents the same function as $\mathscr{A}$. ∎
@@
In Section 3.1 we described a Bayesian network $p_{G}=p^{*}$ with the required graph. It remains to show that this network represents the same distribution as $\mathscr{A}$. We will do this by showing that the Bayesian network has the same distribution as the augmented PC, i.e. $p_{G}(\bm{X},\bm{Z})=p_{\mathscr{A}_{\text{aug}}}(\bm{X},\bm{Z})$.
@@
$p_{\mathscr{A}_{\text{aug}},T}(\bm{X},\bm{Z}):=\prod_{(t_{i},t_{j})\in\textsc{sumedges}(T)}w_{t_{i},t_{j}}\prod_{t\in\textsc{leaves}_{\bm{X}}(T)}f_{t}(X_{\text{sc}(t)})\prod_{t\in\textsc{leaves}_{\bm{Z}}(T)}f_{t}(Z_{\text{sc}(t)})$ (1)
where $\textsc{sumedges}(T)$ denotes the set of outgoing edges from sum nodes in $T$, and $\textsc{leaves}_{\bm{X}}(T),\textsc{leaves}_{\bm{Z}}(T)$ denote the set of leaf nodes in $T$ with scope corresponding to a variable in $\bm{X},\bm{Z}$ respectively. The distribution of the augmented PC is then in fact given by the sum of these functions over all induced trees:
@@
$p_{\mathscr{A}_{\text{aug}}}(\bm{X},\bm{Z})=\sum_{T\in\mathcal{T}}p_{\mathscr{A}_{\text{aug}},T}(\bm{X},\bm{Z})$.
Now, let $\text{path}(v,i,j)$ be a predicate indicating whether there is a path between $t_{p,j}$ and $t_{v,i}$ (where we use $p$ to denote the parent vtree node of $v$, and as before e.g. $t_{v,i}$ indicates the product node with scope $\bm{X}_{v}$ and corresponding to $Z_{v}=i$). We will consider two cases depending on the value of the latents. Specifically, we will say that an assignment $\bm{z}$ is consistent if $\text{path}(v,z_{v},z_{p})$ holds for all non-root inner nodes in the vtree, and inconsistent otherwise.
If an assignment $\bm{z}$ is inconsistent, then for any assignment $\bm{x}$ of the observed variables, we have that $p_{G}(\bm{x},\bm{z})=0$ by definition of the Bayesian network distribution. Now consider any induced subtree $T\in\mathcal{T}$. Each $T$ must contain one product node for every variable scope. In particular, $T$ must contain some product node $t_{v,i}$ such that $z_{v}\neq i$ (otherwise, since $\bm{z}$ is inconsistent, a (connected) tree would be impossible). We then have $p_{\mathscr{A}_{\text{aug}},T}(\bm{x},\bm{z})=0$ for all $\bm{x}$, as Equation 1 then contains a leaf function $f_{t}(z_{v})=\mathds{1}_{z_{v}=i}=0$. Thus $p_{\mathscr{A}_{\text{aug}}}(\bm{x},\bm{z})=\sum_{T\in\mathcal{T}}p_{\mathscr{A}_{\text{aug}},T}(\bm{x},\bm{z})=0$ for any $\bm{x}$.
If an assignment $\bm{z}$ is consistent, note that, by our assumption of alternating sums and products, there can be exactly one path from $t_{p,j}$ to $t_{v,j}$, as $t_{p,j}$ has a unique sum node child with scope containing $\bm{X}_{v}$, and this sum node must immediately have $t_{v,j}$ as a child. Thus there is exactly one induced tree $T$ containing $t_{v,z_{v}}$ for all (non-root) inner vtree nodes $v$. Further, examining the definition of the Bayesian network distribution $p_{G}(\bm{X},\bm{z})$, this exactly matches the definition of $p_{\mathscr{A}_{\text{aug}},T}(\bm{X},\bm{z})$: each sum node edge weight in the tree corresponds to a sum node edge weight along a path from some $t_{p,z_{p}}$ to $t_{v,z_{v}}$ and thus the CPT of $Z_{v}$ given $Z_{p}$ (the root sum node edge weight corresponds to the CPT for $Z_{\textsc{root}(V)}$), and each leaf node distribution for observed variables corresponds to the CPT for that variable given its parent.
@@
We first show that $\bm{C}_{w}$ satisfies the following properties:
@@
    $X_{a:b}=\bigcup_{Z_{i}\in\bm{C}_{a:b}}\textsc{Leaves}(Z_{i})$ is a disjoint union.
@@
    If $a\!\leq\!c\!\leq d\!\leq b$, then for $Z\!\in\!\bm{C}_{c:d}$, there exists $Z^{\prime}\in\bm{C}_{a:b}$ such that $Z^{\prime}$ is an ancestor of $Z$ in $G_{\mathscr{A}}$.
Property 1 follows from the proof of correctness of the segment tree querying algorithm. Property 2 follows from Property 1 together with the key observation that we can compute $\bm{C}_{c:d}$ via $\bigcup_{Z_{i}\in\bm{C}_{a:b}}\textsc{SegmentCover}(Z_{i},\bm{X}_{c:d}\cap\textsc{Leaves}(Z_{i}))$. Let $w=a\!:\!b$ be a node in $W$ with children $l=a\!:\!c$ and $r=c+1\!:\!b$; it follows from Property 1 that $\bm{C}_{a:b}$ covers $\bm{X}_{a:b}$ and $\bm{C}_{a:c}$ blocks all paths from $\bm{X}_{a:c}$ to $\bm{C}_{c+1:b}$; it follows from Property 2 that $\bm{C}_{a:c}$ blocks all paths from $\bm{X}_{a:c}$ to $\bm{X}_{a:b}$. Hence we conclude that $\bm{C}_{w}$ is a valid labelling. A minor catch is that $\bm{C}_{w}$ may contain variables in $\bm{X}$, but we can replace them by their parent in $G_{\mathscr{A}}$ without affecting the validity of $\bm{C}_{w}$.
Now we show that $|\bm{C}_{w}|\leq 4d$ and it suffices to show that for each layer of the vtree $V$, the number of nodes $v$ visited by Algorithm 3 is at most $4$. We prove this by a top-down induction. For the base case, the root node is the only node in the first layer hence the number of nodes visited is 1. Now assume that for the $k$-th layer, the number of nodes visited is at most $4$, and denote them by $v_{1}$, $v_{2}$, $v_{3}$, $v_{4}$, ordered from “left” to “right”. Note that $\{v_{i}\}_{1\leq i\leq 4}$ must form a contiguous segment with $v_{2}$ and $v_{3}$ completely covered by $\bm{X}_{a:b}$; hence the function call on $v_{2}$ and $v_{3}$ would return without further recursive calls. For the function calls on $v_{1}$ and $v_{2}$, suppose both of them trigger recursive calls, in the worse case scenario, each would trigger at most two recursive calls; hence, the number of nodes visited in the $(k+1)$th layer is at most $4$. Since there are $d$ layers in total, the total number of nodes returned is at most $4d$. ∎
@@
By Theorem 5.1, given a structured PC $\bm{X}$ over $n$ variables with hidden state size $h$ and respecting vtree $V$, we can generate a vtree $W$ of depth $O(\log n)$ and with labelling cardinality $M^{\prime}=3$. Thus, by Theorem 3.8 we can construct a PC representing the same function and respecting vtree $W$ of size $O(nh^{3})$. The depth of the PC is then also $O(\log n)$ as we have assumed alternating sum and product nodes, so the depth of the circuit is at most double that of the vtree. ∎
@@ Appendix B Computing Minimum D-separators
procedure MinimumSeparator($Z$, $\bm{A}$, $\bm{B}$)
     if $\bm{A}\!=\!\varnothing$ and $\bm{B}\!=\!\varnothing$ then
@@
           $Z_{i},\bm{A}\!\cap\!\textsc{Leaves}(Z_{i}),\bm{B}\!\cap\!\textsc{Leaves}(Z_{i}))$
@@
     $\bm{C}_{\bm{A}}\leftarrow\textsc{Min}({\bigcup}_{i}\bm{C}_{i}\cup\{Z\},{\bigcup}_{i}\bm{C}_{i,\bm{A}})$
     $\bm{C}_{\bm{B}}\leftarrow\textsc{Min}({\bigcup}_{i}\bm{C}_{i}\cup\{Z\},{\bigcup}_{i}\bm{C}_{i,\bm{B}})$
     $\bm{C}\leftarrow\textsc{Min}(\bm{C_{A}},\bm{C_{B}})$
     return $\bm{C_{A}},\bm{C_{B}},\bm{C}$
@@
Let $G$ be a tree-shaped Bayesian network rooted at $Z$; in particular, assume that the leaves of $G$ $\subseteq\bm{X}\!\cup\!\bm{Z}$ and the internal nodes of $G$ $\subseteq\bm{Z}$ (e.g. Figure 1(b)). Then, we want to prove that Algorithm 5 computes a minimum d-separator $\bm{C}\subseteq G$ for $\bm{A},\bm{B}\subseteq\bm{X}$.
As shown in Algorithm 5, given tree-shaped Bayesian network $G$ rooted at $Z$, the procedure MinimumSepartor computes three sets of latent variables $\bm{C_{A}}$, $\bm{C_{B}}$ and $\bm{C}$. Specifically, we shall prove the following properties:
@@
    $\bm{\bm{C_{A}}}$ is a minimum d-separator between $\bm{A}$ and $\bm{B}$ that also blocks all paths from $\bm{A}$ to $Z$.
@@
    $\bm{\bm{C_{B}}}$ is a minimum d-separator between $\bm{A}$ and $\bm{B}$ that also blocks all paths from $\bm{B}$ to $Z$.
@@
    Either $\bm{C_{A}}$ or $\bm{C_{B}}$ is a minimum d-separator between $\bm{A}$ and $\bm{B}$ in $G$ (rooted at $Z$); hence $\bm{C}$ is a minimum d-separator between $\bm{A}$ and $\bm{B}$ in $G$.
@@
We prove Property 1 (and Property 2) by a bottom-up induction on $G$. First of all it is easy to verify that the three base cases, i.e. $\bm{A}=\varnothing$ and $\bm{B}=\varnothing$, $\bm{A}=\varnothing$ and $\bm{B}=\varnothing$, are correct.
We now prove Property 1 (the proof for Property 2 is symmetric) by induction; first of all it is clear that both ${\bigcup}_{i}\bm{C}_{i}\cup\{Z\}$ and ${\bigcup}_{i}\bm{C}_{i,\bm{A}}$ form d-separators between $\bm{A}$ and $\bm{B}$, and it remains to show that the minimum of these two is a minimum d-separator between $\bm{A}$ and $\bm{B}$ that blocks all paths from $\bm{A}$ to $Z$. Suppose, towards a contradiction, let $\bm{C_{A}}^{\prime}$ be such a d-separator of size $<\textsc{Min}(|{\bigcup}_{i}\bm{C}_{i}\cup\{Z\}|,|{\bigcup}_{i}\bm{C}_{i,\bm{A}}|)$. There are two cases:
  • If $\bm{C_{A}}^{\prime}$ contains $Z$: let $G_{i}$ be the subtree rooted at $Z_{i}$ and set ${\bm{C}_{i}}^{\prime}=\bm{C_{A}}^{\prime}\cap G_{i}$. It is immediate that $\bm{C_{i}}^{\prime}$ is a d-separator between $\bm{A}$ and $\bm{B}$ in $G_{i}$ and by assumption, there exists at least one one $i$ such that $|{\bm{C}_{i}}^{\prime}|<|\bm{C}_{i}|$; contradicting the induction hypothesis.
  • If $\bm{C_{A}}^{\prime}$ does not contain $Z$: let $G_{i}$ be the subtree rooted at $Z_{i}$ and set ${\bm{C}_{i,\bm{A}}}^{\prime}=\bm{C_{A}}^{\prime}\cap G_{i}$. Similarly, it is not hard to see that ${\bm{C}_{i,\bm{A}}}^{\prime}$ is a d-separator between $\bm{A}$ and $\bm{B}$ in $G_{i}$ that blocks all paths from $\bm{A}\cap G_{i}$ to $Z_{i}$. By assumption, there exists at least one $i$ such that $|{\bm{C}_{i,\bm{A}}}^{\prime}|<|{\bm{C}_{i,\bm{A}}}|$; contradicting the induction hypothesis.